OpenStudy (anonymous):

A dealer wishes to produce 300 gallons of oil worth 40 cents a quart. He has oil worth 36 cents a quart and oil worth 52 cants a quart. How many quarts of each should he mix?

OpenStudy (anonymous):

First of all 300 gallons = 1200 quarts. x + y = 1200 .36x + .52y = .40(1200) x + y = 1200 36x + 52y = 40(1200) -36x - 36y = -43200 36x + 52y = 48000 16y = 4800 y = 300 x = 900 So 900 quarts of the 36 cent 300 quarts of the 52 cent